OpenStudy (anonymous):

Solve by factoring x^2-7x=-12

hero (hero):

Add 12 to both sides first before factoring

hero (hero):

x^2 - 7x + 12 = 0

hero (hero):

Find factors of 12 that add up to - 7

hero (hero):

-4 + -3 = -7 (-4)(-3) = 12

hero (hero):

Find the factors of x^2: x*x Write the factors of x^2 -7x + 12 as the product of two binomials: (x-4)(x-3) = 0

hero (hero):

Use Zero Product Property ab = 0 to set each binomial equal to zero: x - 4 = 0 x - 3 = 0

hero (hero):

Solve for x: x = 4 x = 3